OpenStudy (anonymous):

a student earned a total of 500 points on the first 6 tests. what score does she need on the next 2 tests to average 85 points?

OpenStudy (lgbasallote):

ahhh i always do this >:))) let x = score on the last quiz = second last average = total/number of quizzes number of quizzes = 6 + 2 = 8 (500 + x + x)/8 = 85 500 + x + x = 680 2x = 680 - 500 2x = 180 x = 90 she needs 90 pts in each quiz..tough luck

OpenStudy (anonymous):

thanks